# Differentially private projection-depth-based medians

### v1

@@ Appendix A Algorithm
Let $\mu$ be the median functional. If Condition 4 holds for $(1/2,\nu,k)$, then there exists a universal constant $c>0$ such that for all $d,n\geq 1$ and $t>0$ it holds that
$\Pr\left(\tilde{S}_{n,k}(\mu)\geq t\right)\lesssim\exp\left(-nr^{2}M^{2}+d\log(cn/d)\right)+\exp\left(-3MC/4k+d\log 9+\log n\right).$
@@
Therefore, applying Lemma C.2 with $q=1/2$ and $k=k$ gives the result.
@@
$\Pr\Big{(}\tilde{S}_{n,k}(\sigma)\geq C/n\Big{)}\lesssim\exp\left(-nr^{2}M^{2}+d\log(cn/d)\right)\\ +\exp\left(-3MC/8k+d\log 9+\log(n(1-2q^{\prime}+1/n))\right).$
@@
Now, by assumption, $2k<n(q-q^{\prime})$, and so, for all $u\in\mathbb{S}^{d-1}$, there exists $q^{\prime}<p_{u}<1-q^{\prime}$ such that $\sigma(\tilde{\nu}_{u})=|\mu(\tilde{\nu}_{u})-\hat{\xi}_{p_{u},u}|$. (This is because $\hat{\xi}_{q,u}$ is at least $2k$ observations away from $\hat{\xi}_{q^{\prime},u}$.) Therefore, on $E_{10}$, $\lfloor(n+1)/2\rfloor$ of the observations lie within $[\mu(\hat{\nu}_{u})-\sigma(\tilde{\nu}_{u})-C/n,\mu(\hat{\nu}_{u})+\sigma(\tilde{\nu}_{u})+C/n]$, which immediately implies that $\sigma(\hat{\nu}_{u})-C/n\leq\sigma(\tilde{\nu}_{u})$.
@@
Therefore, applying Lemma C.2 with $q=1/4$ and $k=k$ yields the result.
@@ Appendix D Proofs and related lemmas from Section 5
which completes the proof.
@@
Using the fact that $W$ is a standard Laplace random variable, for all $k\in\{0\}\cup[n-1]$, it holds that
@@ Appendix E Results concerning Condition 2
Here, we show that the median and median absolute deviation satisfy Condition 2 under general conditions.
@@
$\inf_{\begin{subarray}{c}0<p<1/2\\ |\xi_{p,u}-\xi_{q,u}|>t\end{subarray}}|\xi_{q,u}-\xi_{p,u}+\xi_{q+1/2,u}-\xi_{p+1/2,u}|=\inf_{\begin{subarray}{c}0<p<1/2\\ |\xi_{p,u}-\xi_{q,u}|>t\end{subarray}}[|\xi_{q,u}-\xi_{p,u}|+|\xi_{q+1/2,u}-\xi_{p+1/2,u}|]\geq t.$
@@
o1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v2

@@ Appendix B Proofs from Section 3
If Condition 4 holds for $(1/2,\nu,k)$, then there exists a universal constant $c>0$ such that for all $d,n\geq 1$ and $t>0$ it holds that
$\Pr\left(\tilde{S}_{n,k}(\operatorname{MED})\geq t\right)\lesssim\exp\left(-nr^{2}M^{2}+d\log(cn/d)\right)+\exp\left(-3MC/4k+d\log 9+\log n\right).$
@@
Therefore, applying Lemma C.2 with $q=1/2$ and $k=k$ gives the result. ∎
@@
$\Pr\Big{(}\tilde{S}_{n,k}(\operatorname{MAD})\geq C/n\Big{)}\lesssim\exp\left(-nr^{2}M^{2}+d\log(cn/d)\right)\\ +\exp\left(-3MC/8k+d\log 9+\log(n(1-2q^{\prime}+1/n))\right).$
@@
Now, by assumption, $2k<n(q-q^{\prime})$, and so, for all $u\in\mathbb{S}^{d-1}$, there exists $q^{\prime}<p_{u}<1-q^{\prime}$ such that $\sigma(\tilde{\nu}_{u})=|\mu(\tilde{\nu}_{u})-\hat{\xi}_{p_{u},u}|$. (This is because $\hat{\xi}_{q,u}$ is at least $2k$ observations away from $\hat{\xi}_{q^{\prime},u}$.) Therefore, on $E_{10}$, $\lfloor(n+1)/2\rfloor$ of the observations lie within $[\mu(\hat{\nu}_{u})-\sigma(\tilde{\nu}_{u})-C/n,\mu(\hat{\nu}_{u})+\sigma(\tilde{\nu}_{u})+C/n]$, which immediately implies that $\sigma(\hat{\nu}_{u})-C/n\leq\sigma(\tilde{\nu}_{u})$. ∎
Proof of Theorem 3.5.
The results follows from Lemma C.3 and Lemma C.4. ∎
@@
Therefore, applying Lemma C.2 with $q=1/4$ and $k=k$ yields the result. ∎
@@ Appendix D Proofs and related lemmas from Section 5
which completes the proof. ∎
@@
By a union bound, for all $k\in\{0\}\cup[n-1]$, it holds that
@@
$\leq\Pr\left(\operatorname{SM}(\hat{\nu},Q)<k\big{|}\frac{2\log(1/2\delta)}{\epsilon}-\frac{2}{\epsilon}W\leq k\right)$
$\hskip 120.00018pt+\Pr\left(\frac{2\log(1/2\delta)}{\epsilon}-\frac{2}{\epsilon}W>k\right)$
@@ Appendix E Results concerning Condition 2
Here, we show that the trimmed estimators, median and median absolute deviation satisfy Condition 2.
@@
$\inf_{\begin{subarray}{c}0<p<1/2\\ |\xi_{p,u}-\xi_{q,u}|>t\end{subarray}}|\xi_{q,u}-\xi_{p,u}+\xi_{q+1/2,u}-\xi_{p+1/2,u}|\\ =\inf_{\begin{subarray}{c}0<p<1/2\\ |\xi_{p,u}-\xi_{q,u}|>t\end{subarray}}[|\xi_{q,u}-\xi_{p,u}|+|\xi_{q+1/2,u}-\xi_{p+1/2,u}|]\geq t.$
@@
Lemma E.3.
Suppose that $\nu=\mathcal{N}(\mu,\Sigma)$. Then there exists a universal constant $c>0$ such that for all $20\leq\alpha<1/$ 2 Condition 2 holds for $(\nu,\operatorname{TM}_{\alpha},\operatorname{TAD}_{\alpha})$ with
$\zeta_{n}(t)=e^{-cn(1-2\alpha)t^{2}/\lVert\Sigma^{-1/2}\rVert^{2}+d}.$
Proof.
WLOG, assume that $\mu=0$, and let $Z_{\alpha}\sim N(0,\Sigma/n(1-2\alpha))$. Then, it follows from Gaussian concentration of measure
$\Pr\left(S_{n}(\operatorname{TM}_{\alpha})\geq t\right)=\Pr\left(\sup_{u}|\bar{X}_{\alpha}^{\top}u|\geq t\right)=\Pr\left(\lVert\bar{X}_{\alpha}\rVert\geq t\right)\lesssim e^{-n(1-2\alpha)t^{2}/\lVert\Sigma^{-1/2}\rVert^{2}}.$ (E.8)
Now, we focus on $\Pr\left(S_{n}(\operatorname{TAD}_{\alpha})\geq t\right)$. First, we have that
$\Pr\left(S_{n}(\operatorname{TAD}_{\alpha})\geq t\right)\leq\Pr\left(S_{n}(\operatorname{TAD}_{\alpha})\geq t/\lVert\Sigma^{-1/2}\rVert\right).$
From now on, assume that $\nu=\mathcal{N}(0,I)$ and let $\mathbb{Z}_{n}\sim N(0,I)$.
$S_{n}(\operatorname{TAD}_{\alpha})$ $=\sup_{u\in\mathbb{S}^{d-1}}|\operatorname{TAD}_{\alpha}(\hat{\nu}_{u})-\operatorname{TAD}_{\alpha}(\nu_{u})|$
$=\sup_{u\in\mathbb{S}^{d-1}}|\frac{1}{n(1-2\alpha)}\sum_{i=\alpha n}^{n-\alpha n}|X_{i}^{\top}u-\bar{X}_{\alpha}^{\top}u|-\int|Z^{\top}u|d\nu|$
$=\sup_{u\in\mathbb{S}^{d-1}}\frac{1}{n(1-2\alpha)}\sum_{i=\alpha n}^{n-\alpha n}||X_{i}^{\top}u-\bar{X}_{\alpha}^{\top}u|-\sqrt{2\pi}|$
$=\sup_{u\in\mathbb{S}^{d-1}}\frac{1}{n(1-2\alpha)}\sum_{i=\alpha n}^{n-\alpha n}||Z_{i}^{\top}u|-\sqrt{2\pi}|+\sup_{u\in\mathbb{S}^{d-1}}|\bar{X}_{\alpha}^{\top}u-\mu_{u}|$
$=\frac{1}{n(1-2\alpha)}\sup_{u\in\mathbb{S}^{d-1}}\lVert|\mathbb{Z}_{n}^{\top}u|-\sqrt{2\pi}\mathbbm{1}\rVert_{1}+\sup_{u\in\mathbb{S}^{d-1}}|\bar{X}_{\alpha}^{\top}u-\mu_{u}|$
$\leq\frac{1}{\sqrt{n(1-2\alpha)}}\sup_{u\in\mathbb{S}^{d-1}}\lVert|\mathbb{Z}_{n}^{\top}u|-\sqrt{2\pi}\mathbbm{1}\rVert_{2}+\sup_{u\in\mathbb{S}^{d-1}}|\bar{X}_{\alpha}^{\top}u-\mu_{u}|$
$\coloneqq I+II.$
Applying this bound, along with a union bound yields that there exists a universal constant $C>0$ such that
$\Pr\left(S_{n}(\operatorname{TAD}_{\alpha})\geq t\right)$ $\leq\Pr\left(I+II\geq Ct\right)\leq\Pr\left(I\geq Ct/2\right)+\Pr\left(II\geq Ct/2\right).$
Now, for $II$, applying (E.8) yields that
$\Pr\left(II\geq Ct/2\right)\leq e^{-n(1-2\alpha)t^{2}/\lVert\Sigma^{-1/2}\rVert^{2}}.$ (E.9)
We now focus on $I$, and apply an $\epsilon$-net argument. Recall that the set $\mathcal{N}_{\epsilon}$ is an $\epsilon$-net on $\mathbb{S}^{d-1}$ if for all $u\in\mathbb{S}^{d-1}$, there exists $v\in\mathcal{N}_{\epsilon}$ such that $\lVert u-v\rVert\leq\epsilon$. Letting $\mathcal{N}_{\epsilon}$ be an arbitrary $\epsilon$-net on $\mathbb{S}^{d-1}$, it is straightforward (see, e.g., [26, Exercise 4.4.4]) to show that for any $c>0$, $y\in\mathbb{R}^{d}$ and $20<\epsilon<1/$ 2,
$\sup_{u\in\mathbb{S}^{d-1}}||y^{\top}u|-c|\leq\frac{C}{1-2\epsilon}\sup_{u\in\mathcal{N}_{\epsilon}}||y^{\top}u|-c|.$
Furthermore, there exists $\mathcal{N}_{1/4}$ such that $|\mathcal{N}_{1/4}|\leq 9^{d}$ [26, see Corollary 4.2.13]. Applying these facts, along with Gaussian concentration of measure yields that there exists a universal constant $c>0$ such that
$\Pr\left(I\geq Ct/2\right)$ $\leq 9^{d}\sup_{u\in\mathcal{N}_{1/4}}\Pr\left(\frac{1}{\sqrt{n(1-2\alpha)}}\lVert|\mathbb{Z}_{n}^{\top}u|-\sqrt{2\pi}\mathbbm{1}\rVert_{2}\right)$
$=9^{d}\Pr\left(\frac{1}{\sqrt{n(1-2\alpha)}}\lVert|\mathbb{Z}_{n}^{\prime}|-\sqrt{2\pi}\mathbbm{1}\rVert_{2}\right)$
$\lesssim e^{-n(1-2\alpha)ct^{2}+d}.$
This completes the proof.
Appendix F Cauchy lower bound
Here, we present a lower bound on the sample complexity of estimating the location parameter from a measure with Cauchy marginals. For a set $S$, let ~$\tilde{\mathcal{M}}_{1,\epsilon,n}(S)$ be the set of maps from $\hat{\nu}$ to $\mathcal{M}_{1}(S)$ which satisfy (2.2). Note that ~$\tilde{\mathcal{M}}_{1,\epsilon,n}(S)$ is essentially the set of all differentially private estimators which lie in $S$. Let $\mathcal{C}$ be the set of measure with Cauchy marginals over $\mathbb{R}^{d}$.
Theorem F.1.
For all $n,d\geq 1$, $t\geq 0$ it holds that
$n=\Omega\left(\frac{d}{t^{2}}\vee\frac{d}{t\epsilon}\right),$
samples are required for
$\inf_{T_{\epsilon}\sim\tilde{\mathcal{M}}_{1,\epsilon,n}(\mathbb{R})}\sup_{\nu\in\mathcal{C}}\mathbb{E}|T_{\epsilon}(\hat{\nu})-\theta(\nu)|\leq t.$ (F.1)
Proof.
[…]
Corollary F.2 ([2]).
For all $\epsilon,\tau>0$ and any $\mathcal{P}\subset\mathcal{M}_{1}(\mathbb{R})$, let $\mathcal{Q}=\{Q_{1},\ldots,Q_{m}\}\subset\mathcal{P}$. If for all $i\neq j$, it holds that $|T_{\epsilon}(Q_{i})-T_{\epsilon}(Q_{j})|\geq 3\tau$, $\operatorname{KL}(Q_{i},Q_{j})\leq\beta$, and $\operatorname{TV}(Q_{i},Q_{j})\leq\gamma$, then then $n^{*}=\Omega\left(\log m(\beta^{-1}\vee(\gamma\epsilon)^{-1})\right)$.
Let $P_{\theta}$ be the Cauchy measure with location parameter $\theta\in\mathbb{R}$ and let $P_{\theta}=\prod_{i=1}^{n}P_{\theta_{i}}$ for $\theta=(\theta_{1},\ldots,\theta_{d})\in\mathbb{R}^{d}$. Now, consider a 1/2-packing of the unit ball in $\mathbb{R}^{d}$: $V=\{v_{1},\ldots,v_{M}\}$. Recall that the packing number of the unit ball is at least $2^{d}$, and so $M\geq 2^{d}$. Next, define $\Theta=\{\theta_{1},\ldots,\theta_{M}\}$, where $\theta_{i}=4\alpha v_{i}$ for some $\alpha>0$. It follows that for $i\neq j$, $\lVert\theta_{i}-\theta_{j}\rVert=4\alpha\lVert v_{i}-v_{j}\rVert\geq\alpha$ and that $\lVert\theta_{i}-\theta_{j}\rVert\leq\lVert\theta_{i}\rVert+\lVert\theta_{j}\rVert\leq 8\alpha$.
Now, let $\operatorname{KL}(P,Q)$ denote the Kullback–Leibler divergence between two measures $P$ and $Q$. We now bound the Kullback–Leibler divergence between $\operatorname{KL}(P_{\theta_{i}},P_{\theta_{j}})$ for $j\neq i$. To this end, we have that
$\operatorname{KL}(P_{\theta_{i}},P_{\theta_{j}})$ $=\sum_{k=1}^{d}\log\left(1+|\theta_{ik}-\theta_{jk}|^{2}/4\right)\leq\lVert\theta_{i}-\theta_{j}\rVert^{2}/4\leq 16\alpha^{2}.$
Applying Corollary F.2, we have that $n^{*}=\Omega(\frac{d}{\delta^{2}}\vee\frac{d}{\delta\epsilon})$. ∎